# Video2BEV: Transforming Drone Videos to BEVs for Video-based Geo-localization

### v3

@@ Abstract
Existing approaches to drone visual geo-localization predominantly adopt the image-based setting, where a single drone-view snapshot is matched with images from other platforms. Such task formulation, however, underutilizes the inherent video output of the drone and is sensitive to occlusions and viewpoint disparity. To address these limitations, we formulate a new video-based drone geo-localization task and propose the Video2BEV paradigm. This paradigm transforms the video into a Bird’s Eye View (BEV), simplifying the subsequent inter-platform matching process. In particular, we employ Gaussian Splatting to reconstruct a 3D scene and obtain the BEV projection. Different from the existing transform methods, e.g., polar transform, our BEVs preserve more fine-grained details without significant distortion. To facilitate the discriminative intra-platform representation learning, our Video2BEV paradigm also incorporates a diffusion-based module for generating hard negative samples. To validate our approach, we introduce UniV, a new video-based geo-localization dataset that extends the image-based University-1652 dataset. UniV features flight paths at $30^{\circ}$ and $45^{\circ}$ elevation angles with increased frame rates of up to 10 frames per second (FPS). Extensive experiments on the UniV dataset show that our Video2BEV paradigm achieves competitive recall rates and outperforms conventional video-based methods. Compared to other competitive methods, our proposed approach exhibits robustness at lower elevations with more occlusions.
[…]
@@ 1 Introduction
Lin et al. [27] G, S 1 image + 1 image Image $45^{\circ}$
Vo et al. [56] G, S 1 image + 1 image Image N/A
Tian et al. [50] G, S 1 image + 1 image Image $45^{\circ}$
@@ 2 Related Work
Image-based Geo-localization. Image-based geo-localization, which is usually regarded as a sub-task of image retrieval, applies image query to determine locations [63, 71]. The primary challenge of this task is the large appearance discrepancy due to different viewpoints across platforms, including ground [17], satellite [11], and drone [9, 76]. Previous methods can be coarsely divided into two families: image-level alignment and feature-level alignment. (1) For image-level alignment, Shi et al. [41] leverage polar transformation to warp satellite images to the ground view. Similarly, Wang et al. [61, 67] transform ground images to the satellite view. Regmi et al. [36] synthesize aerial images from ground images to facilitate matching with satellite view via Generative Adversarial Networks (GANs). Tian et al. [49] employ GANs to transform drone-view images into satellite-view images. Andrea et al. [54] convert drone-view images to ground views through 3D reconstruction but the output is with distortions. (2) For feature-level alignment, Dai et al. [6] and Wang et al. [58] establish feature alignment in a region-correspondence manner. Some methods [26, 46] focus on key-point alignment. Other methods focus on improving the discriminative ability of neural networks with tailored modules, such as lite-transformer encoder [62], layer-to-layer attention block [79, 65], adaptive integration module [47], strong backbones [66], adaptation information consistency module [23], spatially-adaptive denormalization [60] and well-designed loss functions, e.g., semantic augmentation loss [70], contrastive loss [71, 31, 22, 8], dynamic weighted decorrelation regularization [59], peer learning [69], instance loss [74] and the optimal transport [43]. Additionally, other methods [37, 28, 42] fuse the extra orientation meta-information from GNSS with extracted features. However, previous methods overlook the viewpoint variation within drone-view videos, treating them in an image-based setting. Our method is among the early attempts to leverage the viewpoint variation of drone-view videos to transform drone-view video to Bird’s-Eye View (BEV), thereby reducing the viewpoint disparity between drone and satellite views.
@@
Video-based Geo-localization While video understanding has been a focus of the computer vision community for decades, the problem remains challenging due to the complexity added by the time dimension and the volume of data. Early works [3, 40, 12] leverage two-stream convolution networks to fuse spatial semantic information with motion information. Subsequently, attention mechanisms [7, 13, 34] have been introduced for long-term video understanding, such as vanilla self-attention [1, 55], shift window [29, 30], masked auto-encoder [14, 51], and local spatiotemporal attention [45]. Recently, large language models [48] have also shown their superiority in video understanding. In visual geo-localization, videos contain more visual information captured through the camera’s trajectory, which can provide more comprehensive information compared to images. Vyas et al. [57] are the first to collect ground-view data in video format and propose a hierarchical approach for processing clips of ground-view videos. Regmi et al. [37] leverage the geo-temporal proximity between the ground-view videos and GNSS locations to extract coherent features from videos. Expanding to a global scale, Kulkarni et al. [21] introduce a large-scale ground-view video dataset for worldwide geo-localization. Different from the ground-view videos, drone-view videos typically contain multi-view and multi-scale information for the target location [32]. In this paper, we collect drone-view data in video format and propose a video-based geo-localization dataset. Rather than adopting the off-the-shelf video backbone, we propose a Video2BEV transformation to leverage the 3D geometric correspondences and enable a straightforward spatial alignment for matching.
@@ 3 The UniV Dataset
Discussion. The contribution to the community. Different from existing datasets [74, 27, 50], the proposed UniV expands the modality from image to video (see Tab. 1), facilitating the development of robust drone visual geo-localization. A single image provides limited information about the corresponding location. When core areas of the location are occluded, single-image queries can not produce reliable matching results (see Fig. 1). In such cases, the video contains both occluded and unoccluded frames. One frame may contain core-area information to complement another frame and together they can provide robust and complete information required for drone visual geo-location. In this way, all re-implemented methods perform better when adopting video data (see Fig. 3). Moreover, the UniV dataset also introduces a new real-world challenge for drone visual geo-localization. The new elevation angle of $30^{\circ}$ is typical in real-world flightsThe United States and the United Kingdom allow drone flights up to 400 feet; China restricts drones up to 120 meters.. The $30^{\circ}$ elevation angle faces more occlusion cases (see Tab. 1), simulating outputs of real-world drone flights.
@@ 4.2 Hard Negative Sample Synthesis
Negative samples play a significant role in discriminative metric learning. Current negative sample mining strategies [8, 31, 71] cannot ensure the quantity and quality of negative samples, as challenging samples are inherently scarce, and the selected negative samples do not necessarily exhibit similar architectural styles or consistent semantic details as the original samples. In order to bypass these drawbacks, we propose to generate diverse BEV representations as hard negative samples through a fine-tuned diffusion model, which is shown in the right part of Fig. 4. After transforming the drone-view video to BEVs $x_{bev}$ via our Video2BEV transformation, we utilize a visual-LLM [16] to generate captions in a multi-round manner for both BEV and satellite-view images. After obtaining captions for the BEVs and satellite images, we fine-tune a stable diffusion network [39] with LoRA [15] to generate diverse synthetic images, during which we freeze the CLIP text encoder [35]. The outputs of this model are negative samples for the subsequent step. Since the transformed BEV and satellite images share the same top-down viewpoint and semantic content, we conduct inference on the same diffusion model to generate negative samples for both BEV and satellite-view images using their corresponding captions. We provide visualizations of original and synthetic images in Fig. 5. Synthetic negative samples exhibit similar semantic contents as the original ones but with different fine-grained information, thus enhancing the discriminative intra-platform representation learning.
@@ 5 Experiment
Implementation Details. Since the captions for the two views (drone and satellite) employ different wording to describe the same location, we synthesize samples based on the text and generate separate negative samples for each view. All input images are resized to 256$\times$256. We train the first stage of the proposed model with the AdamW optimizer, with a batch size of 140, for 140 epochs, and a learning rate of $2e^{-5}$ and $2e^{-4}$ for the encoder and other modules in the first stage respectively. Then we freeze parameters in the first stage and train the second stage from scratch with a similar training configuration. During the test stage, we utilize the similarity scores from the first stage of ours to select the top 32 samples from the gallery, and then apply the second stage of the proposed model to re-rank these top 32 samples. The framework is implemented with PyTorch.
@@
Table 3: Ablation studies on: (a) Video2BEV transformation, the second stage of our method, and synthetic negative samples. (b) Different training strategies. Train Together: we fine-tune the first stage based on the weights pre-trained on ImageNet [38], and train the second stage from scratch at the same time. Fine-tune: we load fine-tuned first-stage weights on UniV, and then train both the first stage and the second stage. Freeze: we load fine-tuned first-stage weights on UniV, then fix the first-stage weights and only train the second stage from scratch. Notably, the Freeze strategy yields the best results. (c) Re-ranking different top-k samples in the second stage of our method. Considering the balance between performance and testing time, we choose to re-rank top-32 samples. D and S denote Drone and Satellite respectively. R@1 is recall at top1. AP (%) is average precision (high is good).
@@ 5.1 Comparisons with Competitive Methods
Quantitative Results. As shown in Tab. 2(a), we compare the proposed method with other competitive methods on the UniV dataset. The performance of our method has surpassed that of other competitive methods [58, 6, 59, 8]. On the $45^{\circ}$ subset, our method achieves gains of 0.30% Recall and 0.58% AP for satellite $\rightarrow$ drone compared to the second-best method. On the $30^{\circ}$ subset, all methods experience a performance drop. As shown in Fig. 6, we highlight some imperfect reconstructed regions by the Video2BEV transformation. The lower elevation of the drone flights raises more occlusions (see Tab. 1), which also compromises our Video2BEV transformation. Compared to the second best method, our method is still robust, receiving improvements of 3.2% AP for drone $\rightarrow$ satellite and 5.8% AP for satellite $\rightarrow$ drone respectively (see Tab. 2(a)). All methods are compared in the video setting, which means we temporally average the outputs of frames in a video from the drone view. For methods with officially released weights (Sample4Geo [8], DWDR [59]), we test these methods on the $45^{\circ}$ test set directly and subsequently retrain and evaluate these methods on the $30^{\circ}$ subsets. For methods without official weights (LPN [58], FSRA [6]), we retrain them on both the $45^{\circ}$ and $30^{\circ}$ subsets to ensure a fair comparison.
Out-of-Distribution (OOD) Scalability. We also evaluate the model trained on the UniV dataset ($45^{\circ}$) on the unseen SUES-200 $45^{\circ}$ test set in an OOD manner. SUES-200 dataset contains dense frames collected in real-world environments, including real-world light, shadow transformations, and disturbances. We observe that our method shows strong OOD potential, surpassing the runner-up method by more than 7% AP for drone $\rightarrow$ satellite (see Tab. 2(b)).
@@
5.2 Ablation Studies and Further Discussion
@@
Effect of Re-ranking Top-K Samples. During the test stage, we select top-k samples from the gallery leveraging the similarity score from the first stage of our method and subsequently re-rank these samples by the second stage. We conduct hyper-parameter experiments with varying values of top-k, and select k $\in$ $\{8,16,32,64,128,256,512\}$ (see  Tab. 3(c)). Re-ranking the top-512 and top-256 samples yields the best performance and re-ranking the top-256, top-128, top-64, and top-32 samples results in a slight performance drop respectively. Re-ranking the top-16 and top-8 samples leads to a further decline in performance. Considering the balance between the performance and the testing time, we re-rank the top 32 samples as default.
@@ 6 Conclusion
  4. 4.
    Ablation study for loss weights.
7 Visualizations
@@ 8 Failure Case Analysis
We provide additional qualitative visualizations of retrieval results, with a particular focus on the failure cases on the UniV (Fig. 12(a)) and SUES-200 (Fig. 12(b)) datasets. In these cases, the proposed method fails to recall the matched image in top-1. We observe that it is challenging because the recalled top-1 image has a very similar pattern to the query image, particularly in terms of the appearance and structure of the geographic target in the two images. In the first case, all recalled images share a similar structure and the predicted scores are relatively high. In the second case, all recalled images have a white rectangular roof. The roof of the ground truth image turns partly gray, which affects the retrieval prediction of our method. In the third case, recalled top-3 results have similar red roofs, making it challenging for the proposed method to accurately retrieve the ground truth building. In the fourth case, the recalled top-1 image has a similar architectural style to the query and the ground truth image is in shadow. Both factors contribute to an inaccurate retrieval result. In the last two cases, satellite-view photo color is not consistent with that of drone-view, resulting in false positive results.

### v4

@@ Abstract
Existing approaches to drone visual geo-localization predominantly adopt the image-based setting, where a single drone-view snapshot is matched with images from other platforms. Such task formulation, however, underutilizes the inherent video output of the drone and is sensitive to occlusions and viewpoint disparity. To address these limitations, we formulate a new video-based drone geo-localization task and propose the Video2BEV paradigm. This paradigm transforms the video into a Bird’s Eye View (BEV), simplifying the subsequent inter-platform matching process. In particular, we employ Gaussian Splatting to reconstruct a 3D scene and obtain the BEV projection. Different from the existing transform methods, e.g., polar transform, our BEVs preserve more fine-grained details without significant distortion. To facilitate the discriminative intra-platform representation learning, our Video2BEV paradigm also incorporates a diffusion-based module for generating hard negative samples. To validate our approach, we introduce UniV, a new video-based geo-localization dataset that extends the image-based University-1652 dataset. UniV features flight paths at $30^{\circ}$ and $45^{\circ}$ elevation angles with increased frame rates of up to 10 frames per second (FPS). Extensive experiments on the UniV dataset show that our Video2BEV paradigm achieves competitive recall rates and outperforms conventional video-based methods. Compared to other competitive methods, our proposed approach exhibits robustness at lower elevations with more occlusions. The code is available at: https://github.com/HaoDot/Video2BEV-Open.
22footnotetext: Corresponding author.
@@ 1 Introduction
Lin et al[28] G, S 1 image + 1 image Image $45^{\circ}$
Vo et al[57] G, S 1 image + 1 image Image N/A
Tian et al[51] G, S 1 image + 1 image Image $45^{\circ}$
@@ 2 Related Work
Image-based Geo-localization. Image-based geo-localization, which is usually regarded as a sub-task of image retrieval, applies image query to determine locations [64, 72]. The primary challenge of this task is the large appearance discrepancy due to different viewpoints across platforms, including ground [18], satellite [11], and drone [9, 77]. Previous methods can be coarsely divided into two families: image-level alignment and feature-level alignment. (1) For image-level alignment, Shi et al[42] leverage polar transformation to warp satellite images to the ground view. Similarly, Wang et al[62, 68] transform ground images to the satellite view. Regmi et al[37] synthesize aerial images from ground images to facilitate matching with satellite view via Generative Adversarial Networks (GANs). Tian et al[50] employ GANs to transform drone-view images into satellite-view images. Andrea et al[55] convert drone-view images to ground views through 3D reconstruction but the output is with distortions. (2) For feature-level alignment, Dai et al[6] and Wang et al[59] establish feature alignment in a region-correspondence manner. Some methods [27, 47] focus on key-point alignment. Other methods aim to improve the discriminative ability of neural networks with tailored modules, such as lite-transformer encoder [63], layer-to-layer attention block [80, 66], adaptive integration module [48], strong backbones [67], adaptation information consistency module [24], spatially-adaptive denormalization [61] and well-designed loss functions, e.g., semantic augmentation loss [71], contrastive loss [72, 32, 23, 8], dynamic weighted decorrelation regularization [60], peer learning [70], instance loss [75] and the optimal transport [44]. Additionally, other methods [38, 29, 43] fuse the extra orientation meta-information from GNSS with extracted features. However, previous methods overlook the viewpoint variation within drone-view videos, thinking in an image-based setting. Our method is among the early attempts to leverage the viewpoint variation of drone-view videos to transform drone-view video to Bird’s-Eye View (BEV), thereby reducing the viewpoint disparity between drone and satellite views.
@@
Video-based Geo-localization While video understanding has been a focus of the computer vision community for decades, the problem remains challenging due to the complexity added by the time dimension and the volume of data. Early works [3, 41, 12] leverage two-stream convolution networks to fuse spatial semantic information with motion information. Subsequently, attention mechanisms [7, 13, 35] have been introduced for long-term video understanding, such as vanilla self-attention [1, 56], shift window [30, 31], masked auto-encoder [15, 52], and local spatiotemporal attention [46]. Recently, large language models [49] have also shown their superiority in video understanding. In visual geo-localization, videos contain more visual information captured through the camera’s trajectory, which can provide more comprehensive information compared to images. Vyas et al[58] are the first to collect ground-view data in video format and propose a hierarchical approach for processing clips of ground-view videos. Regmi et al[38] leverage the geo-temporal proximity between the ground-view videos and GNSS locations to extract coherent features from videos. Expanding to a global scale, Kulkarni et al[22] introduce a large-scale ground-view video dataset for worldwide geo-localization. Different from the ground-view videos, drone-view videos typically contain multi-view and multi-scale information for the target location [33]. In this paper, we collect drone-view data in video format and propose a video-based geo-localization dataset. Rather than adopting the off-the-shelf video backbone, we propose a Video2BEV transformation to leverage the 3D geometric correspondences and enable a straightforward spatial alignment for matching.
@@ 3 The UniV Dataset
Discussion. The contribution to the community. Different from existing datasets [75, 28, 51], the proposed UniV expands the modality from image to video (see Tab. 1), facilitating the development of robust drone visual geo-localization. A single image provides limited information about the corresponding location. When core areas of the location are occluded, single-image queries can not produce reliable matching results (see Fig. 1). In such cases, the video contains both occluded and unoccluded frames. One frame may contain core-area information to complement another frame and together they can provide robust and complete information required for drone visual geo-location. In this way, all re-implemented methods perform better when adopting video data (see Fig. 3). Moreover, the UniV dataset also introduces a new real-world challenge for drone visual geo-localization. The new elevation angle of $30^{\circ}$ is typical in real-world flights***The United States and the United Kingdom allow drone flights up to 400 feet; China restricts drones up to 120 meters.. The $30^{\circ}$ elevation angle faces more occlusion cases (see Tab. 1), simulating outputs of real-world drone flights.
@@ 4.2 Hard Negative Sample Synthesis
Negative samples play a significant role in discriminative metric learning. Current negative sample mining strategies [8, 32, 72] cannot ensure the quantity and quality of negative samples, as challenging samples are inherently scarce, and the selected negative samples do not necessarily exhibit similar architectural styles or consistent semantic details as the original samples. In order to bypass these drawbacks, we propose to generate diverse BEV representations as hard negative samples through a fine-tuned diffusion model, which is shown in the right of Fig. 4. After transforming the drone-view video to BEVs $x_{bev}$ via our Video2BEV transformation, we utilize a visual-LLM [17] to generate captions for both BEV and satellite-view images. After obtaining captions for the BEVs and satellite images, we fine-tune a stable diffusion network [40] with LoRA [16] to generate diverse synthetic images, during which we freeze the CLIP text encoder [36]. The outputs of this model are negative samples for the subsequent step. Since the transformed BEV and satellite images share the same top-down viewpoint and semantic content, we conduct inference on the same diffusion model to generate negative samples for both BEV and satellite-view images using corresponding captions. We provide visualizations of original and synthetic images in Fig. 5. Synthetic negative samples exhibit similar semantic contents as the original ones but with different fine-grained information, thus enhancing the discriminative intra-platform representation learning.
@@ 5 Experiment
Implementation Details. Since the captions for the two views (drone and satellite) employ different wording to describe the same location, we synthesize samples based on the text and generate separate negative samples for each view. All input images are resized to 256$\times$256. We train the first stage of the proposed model with the AdamW optimizer, with a batch size of 140, for 140 epochs, and a learning rate of $2e^{-5}$ and $2e^{-4}$ for the encoder and other modules in the first stage respectively. Then we freeze parameters in the first stage and train the second stage from scratch with a similar training configuration. During the test stage, we utilize the similarity scores from the first stage to select the top 32 samples from the gallery, and then re-rank these top 32 samples in the second stage. More details are provided in the supplementary materials.
@@
Table 3: Ablation studies on: (a) Video2BEV transformation, the second stage of our method, and synthetic negative samples. (b) Different training strategies. Train Together: we fine-tune the first stage based on the weights pre-trained on ImageNet [39], and train the second stage from scratch. Fine-tune: we load fine-tuned first-stage weights on UniV, and then train both the first stage and the second stage. Freeze: we load fine-tuned first-stage weights on UniV, then fix the first-stage weights and only train the second stage from scratch. Notably, the Freeze strategy yields the best results. (c) Re-ranking different top-k samples in the second stage of our method. Considering the balance between performance and testing time, we choose to re-rank top-32 samples. D and S denote Drone and Satellite, respectively.
@@ 5.1 Comparisons with Competitive Methods
Quantitative Results. As shown in Tab. 2(a), we compare the proposed method with other competitive methods on the UniV dataset. The performance of our method has surpassed that of other competitive methods [59, 6, 60, 8]. On the $45^{\circ}$ subset, our method achieves gains of 0.30% Recall and 0.58% AP for satellite $\rightarrow$ drone compared to the second-best method. On the $30^{\circ}$ subset, all methods experience a performance drop. As shown in Fig. 6, we highlight some imperfect reconstructed regions by the Video2BEV transformation. The lower elevation of the drone flights raises more occlusions (see Tab. 1), which also compromises our Video2BEV transformation. Compared to the second best method, our method is still robust, receiving improvements of 3.2% AP for drone $\rightarrow$ satellite and 5.8% AP for satellite $\rightarrow$ drone, respectively (see Tab. 2(a)). All methods are compared in the video setting, which means we temporally average the outputs of frames in a video from the drone view. For methods with officially released weights (Sample4Geo [8], DWDR [60]), we test these methods on the $45^{\circ}$ test set directly and subsequently retrain and evaluate these methods on the $30^{\circ}$ subsets. For methods without official weights (LPN [59], FSRA [6]), we retrain them on both the $45^{\circ}$ and $30^{\circ}$ subsets to ensure a fair comparison.
Out-of-Distribution (OOD) Scalability. We also evaluate the model trained on the UniV dataset ($45^{\circ}$) on the unseen SUES-200 $45^{\circ}$ test set in an OOD manner. SUES-200 dataset contains dense frames collected in real-world environments, including real-world light, shadow transformations, and disturbances. We observe that our method shows strong OOD potential, surpassing the runner-up method by more than 7% AP for drone $\rightarrow$ satellite (see Tab. 2(b)). More results on robustness against weather and other real-world variants can be found in the Suppl.
@@
5.2 Ablation Study and Further Discussion
@@
[…]
@@ 6 Conclusion
7 Acknowledgment
We acknowledge support from Guangdong Basic and Applied Basic Research Foundation 2025A1515012281, Nanjing Municipal Science and Technology Bureau 202401035, University of Macau MYRG-GRG2024-00077-FST-UMDF, and National Key R&D Program of China (2022ZD0115502), National Natural Science Foundation of China (NO.62461160308, U23B2010), “Pioneer” and “Leading Goose” R&D Program of Zhejiang (No. 2024C01161).
@@ 7 Acknowledgment
  5. 5.
    Inference efficiency.
  6. 6.
    Additional ablation study:
    • ablation study for loss weights;
    • ablation study for FPS;
    • visualizations of different trajectory lengths.
8 Visualization
@@ 8.3 Visualizations of Synthetic Negative Samples
Figure 12: Visualizations of (a) video-based geo-localization under rainy weather setting and (b) comparisons of rainy data augmentations.
9 Out-of-Distribution (OOD) scalability test in rainy weather
Here, we propose a new research direction, namely, video-based geo-localization under rainy weather. Specifically, this setting aims to match drone-view videos (in rainy weather) with geo-tagged satellite images (in clean weather). See Fig. 12a for more details. We perform rainy data augmentation and out-of-distribution (OOD) testing on the $45^{\circ}$ 2 fps test set (see Tab. 4). Compared to Tab. 2a, all methods, including ours, show suboptimal performance, indicating that this setting is a challenging one. We explore two types of combinations of rainy data augmentation and the Video2BEV transformation in our method. The first type applies rainy data augmentation to the clean drone-view videos (Ours in Tab. 4), meaning the augmentation is applied before the Video2BEV transformation. The second type applies rainy data augmentation to the clean BEV videos (Ours in Tab. 4), meaning the augmentation is applied after the Video2BEV transformation. The performance of these two combinations varies, which we attribute to the sensitivity of the vanilla version of 3DGS to the input in the Video2BEV transformation. Predicting accurate camera poses and point clouds from rainy videos remains challenging, leading to the imperfect visual quality of BEVs from rainy inputs (see Fig. 12b). This finally hinders accurate matching. However, we are confident that future advances in efficient 3DGS may help solve this problem. We leave this as a future research direction.
Table 4: Comparisons in terms of an out-of-distribution testing on the rainy weather setting ($45^{\circ}$ test set). $\dagger$ and * denote rainy data augmentation on drone-view videos and BEV-view videos, respectively.
[…]
R@1 AP R@1 AP
LPN 2.28 3.58 24.82 29.64
FSRA 16.83 21.18 44.08 49.42
DWDR 33.67 38.91 55.63 60.85
Sample4Geo 60.48 64.62 72.32 75.71
Ours 30.10 33.75 36.80 40.98
Ours 65.05 67.49 83.02 84.83
@@ 10 Failure Case Analysis
We provide additional qualitative visualizations of retrieval results, with a particular focus on the failure cases on the UniV (Fig. 13(a)) and SUES-200 (Fig. 13(b)) datasets. In these cases, the proposed method fails to recall the matched image in top-1. We observe that it is challenging because the recalled top-1 image has a very similar pattern to the query image, particularly in terms of the appearance and structure of the geographic target in the two images. In the first case, all recalled images share a similar structure, and the predicted scores are relatively high. In the second case, all recalled images have a white rectangular roof. The roof of the ground truth image turns partly gray, which affects the retrieval prediction of our method. In the third case, recalled top-3 results have similar red roofs, making it challenging for the proposed method to accurately retrieve the ground truth building. In the fourth case, the recalled top-1 image has a similar architectural style to the query, and the ground truth image is in shadow. Both factors contribute to an inaccurate retrieval result. In the last two cases, satellite-view photo color is not consistent with that of drone-view, resulting in false positive results.